Prev Next

SW&IT Meeting Minutes 2020-10-12

Attendees: Carl, David, Igor, Matthieu, Stephen, Trey, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

News from management

Previous Action Items

Owner Action Due Date Completion Date

Open topics

Team reports

MB

  • Last week
    • OC, Management
    • ATT mostly
    • Argos-calunit project description
    • review different ITs
  • This week
    • ROWG meeting
    • ATTSW, ATT Core meetings

DE

  • Syno tech support work postponed after Friday's split-brain freakout
  • Wolfgang satisfied with Zoom solution
  • Yang VM upgrades
  • VMWare contract sent to purchasing
  • OSCO doc/shared folder sorted
This week: vacation

TB

  • Last week
    • Getting the LBTO-DMS set up on my VM
    • First steps w/ Plotly
    • Final additions to report
  • This week
    • More experimenting with Plotly
    • Chart studio

YZ

  • Previous:
    • VM upgrades and some setup (Thanks David & Stephen)
    • Some remaining LBC work (needs to setup lbcview/RB_Science)
    • TMS meeting
    • AO-Cravat meeting and work.
  • Next:
    • Some LBC work to continue.
    • Mainly AO-Cravat work.
    • Make some progress on TMS as well.

CH

  • Last week:
    • AGW:
      • Moved UMAC mockup test equipment to my office.
      • Assisted JP and JM with AGW3 (PEPSI) issues.
        • JP: system wasn’t reacting properly to axis commands. Re-uploaded firmware to eliminate possibility of firmware changes. Further testing suggested electronics problem.
        • JM: system was returning errors when attempting azcam image captures. Retried and found image files. Error message is likely spurious, as with non-PEPSI AGWs.
    • ARGOS: Uploaded one VirtualBox vdi file from argos_dx_lgsw to Google Drive, only 53 minutes for 29GB file.
    • DIMM: System/software walk through with SH, discussed potential software changes, including db replacement, update from qt3 to 5, refactor code to make either proper client/server or monolithic.
    • LBC: 64 bit transition: building fits file creator/tester.
  • This week:
    • AGW: test potential PMAC code changes to support new encoder, directly via Pwen, and via firmware upload from OAC simulator.
    • DMS: Record desired Monnit sensor settings and process for updating premier license in wiki.

IS

Last week:
  • OT
    • Strange issue with obs plots window not coming up. Andrew reported on a Windows system. I couldn't find any issues with it and it works in other Windows OS. I asked him to reinstall the OT but it's more of a pain than I thought it will be as they have to manually export and reimport each program. So I added that to the list of OT enhancements.
  • PIT
    • Worked on a number of issues uncovered on AZ close for proposal deadline. Specifically with the resubmission. Although it worked as intended from the users perspective on the backend I added the new reference to the old one. That created some side effects. A new XSLT and scripts had to be patch and data cleaned up.
    • Worked a little bit on the admit pit end but not much progress.
  • LUCI
    • We looked at issue #8205, increasing RunAO timeout. In reality it is just a IIF command timeout but it seems that waiting for AO tends to trigger it.
    • Also #7602 This was an old issue that Olga thought wasn't fixed related to Non-Sidereal offset calculations but there are discrepancies. It might be correct after all I'm following it up.
    • RTD issue similar to #7952, where image selection in drop-downs changed as one changed tabs. This was a fix done a while back. I'm following it up too with Olga & co.

This week:
  • Safety Training
  • PIT
    • the Close for Proposals opening and closing dates, and listing of current proposals per partner to the pit admin tool
  • LUCI
    • Follow up on last week's issues.
    • Add more debugging information in Journalizer DB storing logic. Trying to figure out why sometimes the mask config file doesn't upload.

SH

  • obs6 and obs5 kickstarting -- Leroy ran PXE
  • VPN
    • LDAP
    • Christian
  • DNS issues resolved for Affiliates
  • Add a couple of external sites to split-tunneling
  • Weatherdisplay

next:

  • MAC address collection
  • LBTI web stuff
  • PEPSI machines on mountain
  • obs1 stuff
Topic revision: r9 - 12 Oct 2020, StephenHooper
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback